What is the perimeter of rectangle WXYZ with coordinates W (3, -4), X(3, -6), Y(-1, -6), and Z (-1, -4)?
Answer by ikleyn(52781)

.

I don't know what can be easier than that.

    The side WX is vertical   (x = 3)  of the length of  2 units = (-4-(-6)).


    The side XY is horizontal  (y = -6)  of the length of  4 units = (3 - (-1)).



    It is just enough to calculate the perimeter = 2 * (2 + 4) = 2*6 = 12  units.    ANSWER

Solved.